OpenStudy (anonymous):

A circle has a radius of 9 ft. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth.

OpenStudy (anonymous):

@jeweaver6895 , i need ur help, plz

OpenStudy (anonymous):

i am stuck, plz help me, i have no idea what it is, plz help

OpenStudy (anonymous):

@jeweaver6895 PPPLLLEEEAAASSSEEE HHHEEELLLPPP @TheSmartOne

OpenStudy (babynini):

The formula for finding the area of a circle is: \[r^2 \times \pi\]

OpenStudy (babynini):

So, in your case it would be \[9^2 \times 3.14 \] :)

OpenStudy (anonymous):

i got 254.34

OpenStudy (anonymous):

is that it??

OpenStudy (babynini):

That's perfect!

OpenStudy (anonymous):

do u mind helping with another, i will tag u

OpenStudy (babynini):

(Remember they want you to round it to the nearest 100, so the answer may be 300? I'm not sure though) I will certainly see if I can help!